Physics- mariam was chased by a dog along a slopy land whose tppography makes the sine of angle of that land to be half(1/2). she runs 30m upward as well as the dog , her mass is 4times less than 5 of that of the dog and the mass of the dog is 20kg. take g=10m/s^2. Calculate tge workdone by mariam and the dog (inclination to the angle)
1
Expert's answer
2020-07-17T08:41:03-0400
"W_d=m_dg(0.5l)\\=0.5(1.25\cdot 20)(10)(30)=3750\ J\\W_m=m_mg(0.5l)\\=0.5( 20)(10)(30)=3000\ J"